40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 310 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  1578   Tue May 12 17:26:56 2009 peteUpdateoplevsetmy oplev quad was bad

Pete, Rob

After looking at some oplev noise spectra in DTT, we discovered that the ETMY quad (serial number 115)  was noisy.  Particularly, in the XX_OUT and XX_IN1 channels, quadrants 2 (by a bit more than an order of magnitude over the ETMX ref) and 4 (by a bit less than an order of mag).  We went out and looked at the signals coming out of the oplev interface board; again, channels 2 and 4 were noise compared to 1 and 3 by about these same amounts.  I popped in the ETMX quad and everything looked fine.  I put the ETMX quad back at ETMX, and popped in Steve's scatterometer quad (serial number 121 or possibly 151, it's not terribly legible), and it looks fine.  We zeroed via the offsets in the control room, and I went out and centered both the ETMX and ETMY quads. 

Attached is a plot.  The reference curves are with the faulty quad (115).  The others are with the 121.

 

Attachment 1: bad_oplev_quad.pdf
bad_oplev_quad.pdf
  1577   Tue May 12 15:22:09 2009 YoichiUpdateLSCArm Finesse

Quote:

It looks as if the measured DARM response is skewed by an extra low pass filter at high frequencies. I don't know why is it so.


One large uncertainty in the above estimate is the cavity pole of X-arm because I simply assumed that the ITMX reflectivity to be the designed value.
I think we can directly measure the X-arm finesse from Alberto's absolute length measurements (i.e. from the width of the resonant peaks in his scans).
By looking at Alberto and Koji's posts (elog:1244 elog:838), it looks like the FWHM of the peaks are around 3kHz. With the FSR ~ 3.8MHz, it gives a finesse of about 1300, which is reasonable.
Alberto, can you check your data and measure the FWHM more precisely ?
Note that we want to measure the FWHM of the peak in the *power* of the beat signal. The beat amplitude is proportional to the electric field *amplitude* of the transmitted auxiliary laser. What we need to get a finesse is the FWHM of the transmitted laser *power*. Thus we need to take the power of the beat signal.
  1576   Tue May 12 01:22:51 2009 YoichiUpdateLSCArm loss
Using the armLoss script (/cvs/cds/caltech/scripts/LSC/armLoss), I measured the round trip loss (RTL) of the arms.

The results are:
XARM: RTL= 171 (+/-2) ppm
YARM: RTL = 181 (+/-2) ppm

To get the results above, I assumed that the transmissivity of the ITMs are the same as the designed value (0.005).
This may not be true though.
  1575   Tue May 12 01:11:55 2009 YoichiUpdateLSCDARM response (DC Readout)
I measured the DARM response with DC readout.

This time, I first measured the open loop transfer function of the X single arm lock.
The open loop gain (Gx) can be represented as a product of the optical gain (Cx), the filter (Fx), and the suspension response (S), i.e. Gx = Cx*Fx*S.
We know Fx because this is the transfer function of the digital filters. Cx can be modeled as a simple cavity pole, but we need to know the finesse to calculate it.
In order to estimate the current finesse of the XARM cavity, I ran the armLoss script, which measures the ratio of the reflected light power between the locked and the unlocked state. Using this ratio and the designed transmissivity of the ITMX (0.005), I estimated the round trip loss in the XARM, which was 170 ppm. From this number, the cavity pole was estimated to be 1608Hz.
Using the measured Gx, the knowledge of Fx and the estimated Cx, I estimated the ETMX suspension response S, which is shown in the first attachment.
Note that this is not a pure suspension response. It includes the effects of the digital system time delay, the anti-imaging and anti-aliasing filters and so on.

Now the DARM open loop gain (Gd) can also be represented as a product of the optical gain (Cd), the filter (Fd) and the suspension response (S).
Since the actuations are applied again to the ETMs and we think ETMX and ETMY are quite similar, we should be able to use the same suspension response as XARM for DARM. Therefore, using the knowledge of the digital filter shape and the measured open loop gain, we can compute the DARM optical gain Cd.
The second attachment shows the estimated DARM response along with an Optickle prediction.
The DARM loop gain was measured with darm_offset_dc = 350. Since we haven't calibrated the DARM signal, I don't know how many meters of offset does this number correspond to. The Optickle prediction was calculated using a 20pm DARM offset. I chose this to make the prediction look similar to the measured one, though they look quite different around the RSE peak. The input power was set to 1.7W in the Optickle model (again this is just my guess).

It looks as if the measured DARM response is skewed by an extra low pass filter at high frequencies. I don't know why is it so.
Attachment 1: SUS_Resp.png
SUS_Resp.png
Attachment 2: DARM_Resp.png
DARM_Resp.png
  1574   Mon May 11 12:25:03 2009 josephb,AlexUpdateComputersfb40m down for patching

The 40m frame builder is currently being patched to be able utilize the full 14 TB of the new raid array (as opposed to being limited to 2 TB).  This process is expected to take several hours, during which the frame builder will be unavailable.

  1573   Mon May 11 11:49:20 2009 steveUpdatePSLMOPA cooling water lines are backwards

Quote:
This is 8 days of 10-minute trend.

DTEC is just the feedback control signal required to keep the NPRO's pump diode at a constant temperature.
Its not the amplifier or the actual NPRO crystal's temperature readout.

There is no TEC for the amplifier. It looks like to me that by opening up the flow to the NPRO some more
we have reduced the flow to the amplifier (which is the one that needs it) and created these temperature
fluctuations.

What we need to do is choke down the needle valve and ream out the NPRO block.




I have measured the "input" line temp at the MOPA box 10 C and the "out" line 8 C

This must be corrected.

However look at the 80 days plot of operation where the head temp variation is nothing new
Attachment 1: htempvar80d.jpg
htempvar80d.jpg
  1572   Sun May 10 13:41:17 2009 steveUpdateVACETMY damping restored, VC1 opened

ETMY damping restored.

Cryo  interlock closed VC1 ~2 days ago. P1 is 6.3 mTorr. Cryo temp 12K stable, reset photoswitch and opened VC1

  1571   Sun May 10 13:34:32 2009 carynUpdatePEMUnplugged Guralp channels

I unplugged Guralp EW1b and Guralp Vert1b and plugged in temp sensors temporarily. Guralp NS1b is still plugged in.

  1570   Sat May 9 15:19:10 2009 ranaUpdatePSLLaser head temperature oscillation
This is 8 days of 10-minute trend.

DTEC is just the feedback control signal required to keep the NPRO's pump diode at a constant temperature.
Its not the amplifier or the actual NPRO crystal's temperature readout.

There is no TEC for the amplifier. It looks like to me that by opening up the flow to the NPRO some more
we have reduced the flow to the amplifier (which is the one that needs it) and created these temperature
fluctuations.

What we need to do is choke down the needle valve and ream out the NPRO block.
Attachment 1: Picture_2.png
Picture_2.png
  1569   Sat May 9 02:20:11 2009 JenneUpdatePSLLaser head temperature oscillation

Quote:
After the laser cooling pipe was unclogged, the laser head temperature has been oscillating in 24h period.
The laser power shows the same oscillation.
Moreover, there is a trend that the temperature is slowly creeping up.
We have to do something to stop this.
Or Rob has to finish his measurements before the laser dies.


How's DTEC doing? I thought DTEC was kind of in charge of dealing with these kinds of things, but after our laser-cooling-"fixing", DTEC has been railed at 0, aka no range.

After glancing at DTEC with Dataviewer along with HTEMP and AMPMON (my internet is too slow to want to post the pic while ssh-ed into nodus), it looks like DTEC is oscillating along with HTEMP in terms of frequency, but perhaps DTEC is running out of range because it is so close to zero? Maybe?
  1568   Sat May 9 00:15:21 2009 YoichiUpdatePSLLaser head temperature oscillation
After the laser cooling pipe was unclogged, the laser head temperature has been oscillating in 24h period.
The laser power shows the same oscillation.
Moreover, there is a trend that the temperature is slowly creeping up.
We have to do something to stop this.
Or Rob has to finish his measurements before the laser dies.
Attachment 1: laser.png
laser.png
  1567   Fri May 8 16:29:53 2009 ranaUpdateComputer Scripts / Programselog and NDS
Looks like the new NDS client worked. Attached is 12 hours of BLRMS.
Attachment 1: Untitled.png
Untitled.png
  1566   Fri May 8 16:03:31 2009 JenneUpdatePEMUpdate on Jenne's Filtering Stuff

To include the plots that I've been working on in some form other than on my computer, here they are:

First is the big surface plot of all the amplitude spectra, taken in 10min intervals on one month of S5 data. The times when the IFO is unlocked are represented by vertical black stripes (white was way too distracting).  For the paper, I need to recreate this plot, with traces only at selected times (once or twice a week) so that it's not so overwhelmingly large.  But it's pretty cool to look at as-is.

Second is the same information, encoded in a pseudo-BLRMS.  (Pseudo on the RMS part - I don't ever actually take the RMS of the spectra, although perhaps I should).  I've split the data from the surface plot into bands (The same set of bands that we use for the DMF stuff, since those seem like reasonable seismic bands), and integrated under the spectra for each band, at each time.  i.e. one power spectra gives me 5 data points for the BLRMS - one in each band.  This lets us see how good the filter is doing at different times.

At the lower frequencies, after ~25 days, the floor starts to pick up.  So perhaps that's about the end of how long we can use a given Wiener filter for.  Maybe we have to recalculate them about every 3 weeks.  That wouldn't be tragic. 

I don't really know what the crazy big peak in the 0.1-0.3Hz plot is (it's the big yellow blob in the surface plot).  It is there for ~2 days, and it seems awfully symmetric about it's local peak.  I have not yet correlated my peaks to high-seismic times in the H1 elog.  Clearly that's on the immediate todo list. 

Also perhaps on the todo list is to indicate in some way (analagous to the black stripes in the surface plot) times when the data in the band-limited plot is just extrapolated, connecting the dots between 2 valid data points.

 

A few other thoughts:  The time chosen for the training of the filter for these plots is 6:40pm-7:40pm PDT on Sept 9, 2007 (which was a Sunday night).  I need to try training the filter on a more seismically-active time, to see if that helps reduce the diurnal oscillations at high frequency.  If that doesn't do it, then perhaps having a "weekday filter" and an "offpeak" filter would be a good idea.  I'll have to investigate.

Attachment 1: H1S5OneMonthWienerCompBLACK.png
H1S5OneMonthWienerCompBLACK.png
Attachment 2: H1S5BandLimitedTimePlot.png
H1S5BandLimitedTimePlot.png
  1565   Fri May 8 15:40:44 2009 peteUpdateLockingprogressively weaker locks

the align script was run after the third lock here.  it would have been interesting to see the arm powers in a 4th lock 

Attachment 1: powers_3lock.pdf
powers_3lock.pdf
  1564   Fri May 8 10:05:40 2009 AlanOmnistructureComputersRestarted backup since fb40m was rebooted

Restarted backup since fb40m was rebooted.

  1563   Fri May 8 04:46:01 2009 rana, yoichiSummaryoplevsBS/PRM/SRM table bad!
We went to center the oplevs because they were far off and found that (as usual) the numbers changed
a little after we carefully centered the oplevs and came back to the control room.

To see if the table was on something soft, we tried pushing the table: no significant effect with ~10 pounds of static force.

With ~10 pounds of vertical force, however, we saw a large change: ~0.25 Oplev units. This corresponds to
~20-30 microradians of apparent optic pitch.

In the time series below you can see the effects:

2.5 s: lid replaced on table after centering.

2.5 - 11 s: various force tests on table

11 s: pre-bias by aligning beams to +0.25 in pitch and then add lid.


So there's some kind of gooey behavior in the table. It takes ~1 s to
settle after we put the lid on. Putting the laptops on the table also
has a similar effect. Please do not put anything on this table lid.
Attachment 1: a.png
a.png
  1562   Fri May 8 04:31:35 2009 ranaUpdateComputer Scripts / Programselog and NDS
In the middle of searching through the elog, its stopped responding. So I followed the Wiki instructions
and restarted it (BTW, don't use the start-elog-nodus script that's in that directory). Seems OK now,
but I am suspicious of how it sometimes does the PDF preview correctly and sometimes not. I found a
'gs' process on there running and taking up > 85% of the CPU.

I also got an email from Chris Wipf at MIT to try out this trick from LASTI to maybe fix the
problems I've been having with the DMF processes failing after a couple hours. I had compiled but
not tested the stuff a couple weeks ago.

Today after it failed, I tried running other stuff in matlab and got some "too many files open" error messages.
So I have now copied the 32-bit linux NDS mex files into the mDV/nds_mexs/ directory. Restarted the
seisBLRMS.m about an hour ago.
  1561   Fri May 8 02:39:02 2009 pete, ranaUpdateLockingcrossover

attached plot shows MC_IN1/MC_IN2.  needs work.

This is supposed to be a measurement of the relative gain of the MCL and AO paths in the CM servo. We expect there to

be a more steep slope (ideally 1/f). Somehow the magnitude is very shallow and so the crossover is not stable. Possible

causes? Saturations in the measurement, broken whitening filters, extremely bad delay in the digital system? needs work.

 

Attachment 1: crossover.pdf
crossover.pdf
Attachment 2: photo.jpg
photo.jpg
  1560   Fri May 8 02:08:59 2009 peteUpdateLockinglock stretches

locks last for about an hour.  this was true last night as well (see "arm power curve" entries).   the second lock shown here evolves differently for unknown reasons.  the jumps in the arm powers of the first lock are due to turning on DC readout.  length-to-angle needs tuning.

 

 

Attachment 1: powers_oplev.pdf
powers_oplev.pdf
  1559   Thu May 7 23:34:59 2009 robUpdateSEIseisBLRMS already lost

Can't find hostname 'fb40m'

 

it only lasted a few hours

  1558   Thu May 7 23:21:04 2009 peteUpdateLockingarm power curve

Quote:

I've plotted TRX, TRY, PD12I and PD11Q.  Arm powers after locking increase for a few tens of minutes, peak out, and then decrease before lock is lost.

 

 

 I should have mentioned that the AS port camera image seems to get progressively uglier over the course of these locks.  Maybe we can use the JoeCam to make a movie of it. 

  1557   Thu May 7 18:12:12 2009 peteUpdateLockingarm power curve

I've plotted TRX, TRY, PD12I and PD11Q.  Arm powers after locking increase for a few tens of minutes, peak out, and then decrease before lock is lost.

 

 

Attachment 1: 2009_may_7_powers.jpg
2009_may_7_powers.jpg
  1556   Thu May 7 17:59:23 2009 AlbertoConfiguration MC WFS
This afternoon the MC could not get locked.
I first checked the Osems values at the MC mirrors and compared them to the trend of the last few hours. That showed that the alignment of the mirrors had slightly changed. I then brought each mirror back to its old alignment state.
 
That let the LSC loop lock the MC, although the reflected power was still high (1.5V) and the WFS control wouldn't engage.
 
Since earlier during the day I was working on the AS table, it is possible that I inadvertently hit the MC REFL beam splitter misaligning the beam to the MC WFS.
To exclude that there was a problem in the suspensions, before touching the WFS, I checked that the cables at the MC's ends and those going to the ADC in the rack were well pushed in.
 
Then I proceeded in centering the beam on both the WFS, balancing the power over the QPDs.
 
In the end the MC could lock again properly.
 
  1555   Thu May 7 15:22:19 2009 josephb, albertoConfigurationComputersfb40m

Quote:

Having determined that Rana (the computer) was having to many issues with testing the new Raid array due to age of the system, we proceeded to test on fb40m.

 

We brought it down and up several times between 11 and noon.  We  eventually were able to daisy chain the old raid and the new raid so that fb40m sees both.  At this time, the RAID arrays are still daisy chained, but the computer is setup to run on just the original raid, while the full 14 TB array is initialized (16 drives, 1 hot spare, RAID level 5 means 14 TB out of the 16 TB are actually available).  We expect this to take a few hours, at which point we will copy the data from the old RAID to the new RAID (which I also expect to take several hours).  In the meantime, operations should not be affected.  If it is, contact one of us.

 

 

 

 

This afternoon the alignment script chrashed after returning sysntax errors. We found that the tpman wasn't running on the framebuilder becasue it had probably failed to get restarted in one of the several reboots executed in the morning by Alex and Jo.

Restarting the tpman was then sufficient for the alignment scripts to get back to work.

  1554   Thu May 7 12:21:36 2009 josephb, alexConfigurationComputersfb40m

Having determined that Rana (the computer) was having to many issues with testing the new Raid array due to age of the system, we proceeded to test on fb40m.

 

We brought it down and up several times between 11 and noon.  We  eventually were able to daisy chain the old raid and the new raid so that fb40m sees both.  At this time, the RAID arrays are still daisy chained, but the computer is setup to run on just the original raid, while the full 14 TB array is initialized (16 drives, 1 hot spare, RAID level 5 means 14 TB out of the 16 TB are actually available).  We expect this to take a few hours, at which point we will copy the data from the old RAID to the new RAID (which I also expect to take several hours).  In the meantime, operations should not be affected.  If it is, contact one of us.

 

 

  1553   Thu May 7 10:28:20 2009 steveUpdateVACretrofitted maglev's needs

 

 Our spare Osaka maglev purchased in Oct 2005 turned out to be having a viton o-ring seal connection on the intake.

It was shipped back to San Jose for retrofitting it with 6" conflat flange ( CF ) This CF is using copper gasket so there will be no permeation of He when you leak check the IFO

 

The digital controller and cable are here. The controller needs to be interfaced with the interlocks and computer system; those have been in a neglected condition lately.

see elog #1505  Historically after every REBOOT of c1vac2 the readbacks works for 3-4 days only. Fixing of this was postponed many times in the past as low priority or lack of knowledgeable

enthusiast.

 

The maglev TG390MCAB wil be back on Tuesday, May 4, 2009.  The mourning of our fateful 360 will only end at the first levitation of the 390.

 

  1552   Wed May 6 19:04:11 2009 ranaSummaryVACvac images
Since there's no documentation on this besides Steve's paper notebooks...

and BTW, since when did the elog start giving us PNG previews of PDFs?
Attachment 1: vacrack.pdf
vacrack.pdf
  1551   Wed May 6 16:56:35 2009 rana, alex, joeConfigurationComputersdaqd log, cront, etc.
While Alex came over, we investigated the log file problems with DAQD and NDS on FB0. There was a lot of
the standard puzzling and mumbling, but eventually we saw that it doesn't create its log file and so it
doesn't write to it. The log file is /usr/controls/main_daqd.log. The other files called daqd.log.DATE
in the logs/ directory are actually not written to. Its awesome.

We also have put in a fix for the overflowing jobs/ directory. It gets a file written to it every time
you make and NDS request and our seisBLRMS has been overloading it. There's now a cron for it in the fb0
crontab which cleans out week-old files at 6:30 AM every day.

We also changed the time of the daily backup from 3:30 AM (when people are still working) to 5:50 AM
(by which time the seismic has ramped up and interferometerists should be asleep). I didn't like the
idea of a bandwidth hog nailing the framebuilder during the peak of interferometer work.

#
# Script to backup via rsync the most recent 40m minute trends and
# any changes to the /cvs/cds filesystem.
#
50 05 * * * /cvs/cds/caltech/scripts/backup/rsync.backup < /dev/null > /cvs/cds/caltech/scripts/\
backup/rsync.backup.log 2>&1

30 06 * * * find /usr/controls/jobs -mtime +7 -exec /bin/rm -f {} \;

seisBLRMS.m restarted on mafalda.
  1550   Wed May 6 02:39:20 2009 YoichiHowToLockingHow to go to DC readout
I wrote a script called DC_readout, which you can find in /cvs/cds/caltech/scripts/DRFPMI/bang/nospring/.

Currently, the locking script succeeds 1/3 of the time. The freaky parts are the MC_F hand off and REFL_DC hand off.
MC_F hand off succeeds 70% of the time. REFL_DC goes well about a half of the time. Combined, the success rate is about 1/3.
We need some work on those hand offs.
Once you pass those freaky parts, the cm_step script usually goes smoothly and you will reach the full RF lock with the boost and the super boost1 engaged on the CM board.

To go to DC readout from there, run the DC_readout script.
First, this script will put some offset to the DARM loop so that some carrier light will leak to the AS port.
You are prompted to lock the OMC. Move the OMC length offset slider to find the carrier resonance and lock the OMC.
You have to make sure that it is carrier, not the 166MHz sideband. Usually the carrier light pulsates around 10Hz or so whereas the 166MHz SB is stable.
Once you locked the OMC to the carrier, hit enter on the terminal running the DC_readout script.
The script will do the rest of the hand off.
Once the script has finished, you may want to check darm_offset_dc in the C1LSC_LA_SET screen. This value sets the DC offset (a.k.a. the homodyne phase).
You may want to change it to what you want.
  1549   Tue May 5 14:02:16 2009 robUpdateLSCDARM DC response varies with DARM offset

Note the effect of quadrature rotation for small offsets.

Attachment 1: DARM_DARM_AS_DC_2.png
DARM_DARM_AS_DC_2.png
Attachment 2: DARM_DARM_AS_DC_3.png
DARM_DARM_AS_DC_3.png
Attachment 3: DARM_DARM_AS_DC_2.pdf
DARM_DARM_AS_DC_2.pdf
Attachment 4: DARM_DARM_AS_DC_3.pdf
DARM_DARM_AS_DC_3.pdf
  1548   Tue May 5 11:44:33 2009 robUpdateLocking DARM response

Here's the RF DARM optical response, on the anti-spring side, from optickle. Note that for the f1 sideband, changing the demod phase mostly adjusts the overall gain, while for the f2 sideband a change in demod phase alters the shape of the response. This is the quadrature-selecting power of using a single RF sideband as a local oscillator.
Attachment 1: DARMtf_nospring.png
DARMtf_nospring.png
Attachment 2: DARMtf_demodphases.png
DARMtf_demodphases.png
  1547   Tue May 5 10:42:18 2009 steveUpdateMOPAlaser power is back

Quote:

As PSL-126MOPA_DTEC went up, the power out put went down yesterday

 The NPRO cooling water was clogged at the needle valve. The heat sink temp was around ~37C

The flow-regulator  needle valve position is locked with a nut and it is frozen. It is not adjustable. However Jeenne's tapping and pushing down on the plastic hardware cleared the way for the water flow.

We have to remember to replace this needle valve when the new NPRO will be swapped in. I checked on the heat sink temp this morning. It is ~18C

There is condensation on the south end of the NPRO body, I wish that the DTEC value would just a little higher like 0.5V

The wavelenght of the diode is temp dependent: 0.3 nm/C. The fine tuning of this diode is done by thermo-electric cooler ( TEC )

To keep the diode precisely tuned to the absorption of the laser gain material the diode temp is held constant using electronic feedback control.

This value is zero now.

 

Attachment 1: uncloged.jpg
uncloged.jpg
  1546   Tue May 5 09:22:46 2009 carynUpdatePEMzeros

For several of the channels on the PEM ADCU, zeros are occuring at the same time. Does anyone know why that might happen or how to fix it?

Attachment 1: zerotest2.png
zerotest2.png
Attachment 2: zerotest.png
zerotest.png
  1545   Tue May 5 08:26:56 2009 robUpdateLockingDC Readout and DARM response

Quote:
Tonight, I was able to switch the DARM to DC readout a couple of times.
But the lock was not as stable as the RF DARM. It lost lock when I tried to measure the DARM loop gain.

I also measured DARM response when DARM is on RF.
The attached plot shows the DARM optical gain (from the mirror displacement to the PD output).
The magnitude is in an arbitrary unit.

I measured a transfer function from DARM excitation to the DARM error signal. Then I corrected it for the DARM open loop gain and the pendulum response to get the plot below.

There is an RSE peak at 4kHz as expected. The origin of the small bump and dip around 2.5kHz and 1.5kHz are unknown.
I will consult with the Optickle model.
I don't know why the optical gain decreases below 50Hz (I don't think it actually decreases).
Seems like the DARM loop gain measured at those frequencies are too low.
I will retry the measurement.


The optical gain does decrease below ~50Hz--that's the optical spring in action. The squiggles are funny. Last time we did this we measured the single arm TFs to compensate for any tough-to-model squiggles in the transfer functions which might arise from electronics or the suspensions.
  1544   Tue May 5 05:16:12 2009 YoichiUpdateLockingDC Readout and DARM response
Tonight, I was able to switch the DARM to DC readout a couple of times.
But the lock was not as stable as the RF DARM. It lost lock when I tried to measure the DARM loop gain.

I also measured DARM response when DARM is on RF.
The attached plot shows the DARM optical gain (from the mirror displacement to the PD output).
The magnitude is in an arbitrary unit.

I measured a transfer function from DARM excitation to the DARM error signal. Then I corrected it for the DARM open loop gain and the pendulum response to get the plot below.

There is an RSE peak at 4kHz as expected. The origin of the small bump and dip around 2.5kHz and 1.5kHz are unknown.
I will consult with the Optickle model.
I don't know why the optical gain decreases below 50Hz (I don't think it actually decreases).
Seems like the DARM loop gain measured at those frequencies are too low.
I will retry the measurement.
Attachment 1: DARM-TF.png
DARM-TF.png
  1543   Mon May 4 16:49:56 2009 AlbertoUpdateMOPAlaser power is dropped

Quote:

As PSL-126MOPA_DTEC went up, the power out put went down yesterday

Alberto, Jenne, Rob, Steve,
 
later on in the afternoon, we realized that the power from the MOPA was not recovering and we decided to hack the chiller's pipe that cools the box.
 
Without unlocking the safety nut on the water valve inside the box, Jenne performed some Voodoo and twisted a bit the screw that opens it with a screw driver. All the sudden some devilish bubbling was heard coming from the pipes.
The exorcism must have freed some Sumerian ghost stuck in our MOPA's chilling pipes (we have strong reasons to believe it might have looked like this) because then the NPRO's radiator started getting cooler.
I also jiggled a bit with the valve while I was trying to unlock the safety nut, but I stopped when I noticed that the nut was stuck to the plastic support it is mounted on.
 
We're now watching the MOPA power's monitor to see if eventually all the tinkering succeeded.

 

[From Jenne:  When we first opened up the MOPA box, the NPRO's cooling fins were HOT.  This is a clear sign of something badbadbad.  They should be COLD to the touch (cooler than room temp).  After jiggling the needle valve, and hearing the water-rushing sounds, the NPRO radiator fins started getting cooler.  After ~10min or so, they were once again cool to the touch.  Good news.  It was a little worrisome however that just after our needle-valve machinations, the DTEC was going down (good), but the HTEMP started to rise again (bad).  It wasn't until after Alberto's tinkering that the HTEMP actually started to go down, and the power started to go up.  This is probably a lot to do with the fact that these temperature things have a fairly long time constant. 

Also, when we first went out to check on things, there was a lot more condensation on the water tubes/connections than I have seen before.  On the outside of the MOPA box, at the metal connectors where the water pipes are connected to the box, there was actually a little puddle, ~1cm diameter, of water. Steve didn't seem concerned, and we dried it off.  It's probably just more humid than usual today, but it might be something to check up on later.]

  1542   Mon May 4 10:38:52 2009 steveUpdateMOPAlaser power is dropped

As PSL-126MOPA_DTEC went up, the power out put went down yesterday

Attachment 1: dtecup.jpg
dtecup.jpg
  1541   Sun May 3 22:48:12 2009 YoichiUpdateLockingSome measurements at the lock point
I attached some measurement results at when the IFO is at the full lock point.

The first plot shows the trend of the arm powers after the interferometer was locked.
The arm powers slowly increased after the lock. This increase is observed every time the IFO is locked.
Probably this is some sort of a thermal effect (mirror lensing, PD efficiency etc).

The second plot is a CARM offset sweep. Even after the demodulation phase optimization, the lock point is not exactly at the resonance.

The third plot is the open loop TF of the AO path. The CM loop UGF is about 20kHz.
The boost and the superboost1 were turned on when this TF was measured. The IFO loses lock if the superboost2 is turned on.

TO DO LIST
Measured the DARM loop shape.
I could not turn on the dewhitening filter for ETMY. ETMX had no trouble. I will check the dewhitening circuit.
Attachment 1: ArmPowerTrend.png
ArmPowerTrend.png
Attachment 2: CARMSweep.png
CARMSweep.png
Attachment 3: CM-AO-Loop-SB1.png
CM-AO-Loop-SB1.png
  1540   Sat May 2 16:34:31 2009 carynDAQPEMGuralp channels plugged back in

I plugged the Guralp cables back into the PEM ADCU

       Guralp NS1b ---> #11

       Guralp Vert1b --->#10

       Guralp EW1b --->#12

  1539   Fri May 1 18:51:34 2009 AlbertoSummaryEnvironmentearthquake

Earthquake 4.4 Leo Carrillo Beach.

Some of the watchdogs tripped out.

  1538   Fri May 1 18:24:36 2009 AlbertoSummaryGeneraljitter of REFL beam ?
Some loud thinking.
 
For the measurement of the length of the PRC, I installed a fast photodiode in the path of the beam reflected by PRM which goes to the 199 PD on the AS table. I picked up the beam by a flipping mirror on the same table.
I have the problem that the DC power that I measure at the PD when the PRC is locked is not constant but fluctuates. This fluctuation is irregular and has a frequency of about one Hz or so. I.e. it makes the 33 Mhz line on the PD oscillate by +/- 10 dBm.
 
Since this fluctuation does not appear at the REFL 33 PD, which has a much larger surface, but also shows up on the REFL 199 PD, I suspected that it was due to the very small size of the fast PDs. If the spot is too large, I thought, the power on the PD should be affected by the beam jitter.
 
Trying to avoid any beam jitter, I placed two lenses with focal lengths one ten times the other on the optical path in such a way to reduce the spot size on my fast PD by the same factors. The DC power was still fluctuating, so I'm not sure it's beam jitter anymore.
 
SPOB is definitely not constant when the PRC is normally locked, even with high loop gains, so maybe the reflected power really fluctuates that much.
Although, if it's actually the DC power that is fluctuating, shouldn't it appear also at the REFL 33 and shouldn't it be a problem that it shows up also in REFL 199? The elog doesn't say anything about that.
 

It's crucial that I get a stable transmitted power to have an accurate measurement of the PRC transmissivity and thus of its macroscopic length.

  1537   Fri May 1 10:04:10 2009 robUpdateLocking166MHz LO phase adjustment

Quote:
I continued to adjust the REFL_2I demodulation phase.
I first optimized the demod phase for SRCL in the DRMI configuration (the error signals were DDs).
Then I restored the full IFO and offset locked it.
Before handing the DARM to RF, I adjusted the 166MHz delay line to maximize the SRCL signal at REFL_2I.
I did this before the DARM RF hand off because changing the delay line setting also changes the AS166 demodulation phase.
After this, I adjusted the digital phase shifter for AS166 to maximize the DARM signal for this port.

I also adjusted the digital demodulation phase of PD11 (REFL_2I) because the optimal demodulation phase for the initial lock acquisition is somewhat (15deg)
different from the optimal demodulation phase for the SRCL when the central part is locked with the DD signals.
This happens because the resonant condition of the central part (lock points of the recycling cavities) changes when the error signals are switched to the DD signals,
due to the offset in the DD signals. This is not good and should be fixed by the optimization of the DD demodulation phases.

Finally, I reduced the CARM offset to zero and tweaked the delay line a bit to maximize the arm power.

Right now, the locking script runs fine until the end.
At the end of the script, I was able to engage the boost on the CM board.



Awesome. Up next: dewhitening.
  1536   Fri May 1 01:32:43 2009 YoichiUpdateLocking166MHz LO phase adjustment
I continued to adjust the REFL_2I demodulation phase.
I first optimized the demod phase for SRCL in the DRMI configuration (the error signals were DDs).
Then I restored the full IFO and offset locked it.
Before handing the DARM to RF, I adjusted the 166MHz delay line to maximize the SRCL signal at REFL_2I.
I did this before the DARM RF hand off because changing the delay line setting also changes the AS166 demodulation phase.
After this, I adjusted the digital phase shifter for AS166 to maximize the DARM signal for this port.

I also adjusted the digital demodulation phase of PD11 (REFL_2I) because the optimal demodulation phase for the initial lock acquisition is somewhat (15deg)
different from the optimal demodulation phase for the SRCL when the central part is locked with the DD signals.
This happens because the resonant condition of the central part (lock points of the recycling cavities) changes when the error signals are switched to the DD signals,
due to the offset in the DD signals. This is not good and should be fixed by the optimization of the DD demodulation phases.

Finally, I reduced the CARM offset to zero and tweaked the delay line a bit to maximize the arm power.

Right now, the locking script runs fine until the end.
At the end of the script, I was able to engage the boost on the CM board.
  1535   Thu Apr 30 15:10:54 2009 robUpdateLockingCARM RF changed to REFL_2I

Quote:
Yoichi, Peter

As Rob suggested, the optimal demodulation phase is easier to find for REFL_2I than POX_1I.
Moreover, for 166MHz LO, we have a phase shifter (delay line) already installed. So we can easily change the demodulation phase of REFL_2I.
Tonight, we switched the RF CARM signal to REFL_2I.
To do so, I changed the signal going to the REFL1 input of the common mode board from POX_1I to REFL_2I.
I moved a BNC-T installed at the output of POX_1I to the REFL_2I output to split the REFL_2I signal and send it to the CM board.
Since the gain of the REFL_2I was about 20dB lower than that of POX_1I, I increased the gain of the SR560, which is installed between the REFL_2 demodulation board and the CM board, from 1 to 10.

With some gain tweaks, we were able to hand off the CARM from REFL_DC to REFL_2I. We also succeeded in switching the REFL_2I ADC channel from PD11 to PD2_DC (the output of the length path from the CM board). This switching is necessary in order to engage the boost on the CM board.

There remains some offset in the CARM when the arm power is maximized. This is expected because the REFL_2I demodulation phase is probably not exactly right.
I will optimize the demodulation phase tomorrow.



From Optickle simulations, it looks like the SRCL/CARM gain ratio at REFL I2 is about 8e-4. So a 1 nanometer offset in SRCL yields 0.8 picometers of offset in CARM.
  1534   Thu Apr 30 05:49:06 2009 YoichiUpdateLocking166MHz LO phase changed
In order to optimize the REFL_2I demod phase, I changed the delay line setting for the 166MHz LO.
Right now, the delay is not yet optimal.
Since the AS166 shares the same LO, the digital demodulation phase of the AS166 had to be changed too.
The DD demod phases and the DD hand off script were also tweaked to improve the resonant condition of the central part.
Now we have more 166MHz coming out of the AS port and the SPOB is larger (more 33MHz resonant in PRC).

Since REFL166 and AS166 demodulation phases are not yet optimized, the cm_step script won't work at this moment.
  1533   Wed Apr 29 15:56:43 2009 robUpdateLockingeffect of SRCL detune on ARM powers in a CARM sweep

With no DARM offset, sweeping CARM shows an asymmetry between the state where we lock to a DARM spring and the state with a DARM anti-spring.  This is why we have a link between the DARM and CARM optical springs. 

For each DARM detune direction (positive or negative, spring or anti-spring), there is only one CARM direction which can yield a DC-based error signal lock with a CARM offset but no DARM offset, which is what we want.

Attachment 1: CARMsweep_DARMspringnospring.png
CARMsweep_DARMspringnospring.png
  1532   Wed Apr 29 10:20:14 2009 steveHowToVACcryo pump interlock rule is waved

Quote:

I tested the cryopump interlock today. It is touchy. I do not have full confidence in it.

I'm proposing that VC1 gate valve should be kept closed while nobody is working in the 40m lab.

How to open gate valve:

1, confirm temp of 12K on the gauge at the  bottom of the cryopump

2, if medm screen cryo reads OFF( meaning warm) hit reset will result reading ON (meaning cold 12K )

3, open VC1 gate valve if P1 is not higher than 20 mTorr

 

VC1 was closed at 18:25,

IFO condition: not pumped,

expected leak plus  out gassing should be less than 5 mTorr/day

The RGA is in bg-mode, annuloses are closed off

 The Cryo pump is running reliably since April 22 hence there is no need to close VC1 repeatedly.

The photo switch interlock was put back onto the H2 vapor pressure gauge and it is working.

  1531   Wed Apr 29 04:03:51 2009 YoichiUpdateLockingCARM RF changed to REFL_2I
Yoichi, Peter

As Rob suggested, the optimal demodulation phase is easier to find for REFL_2I than POX_1I.
Moreover, for 166MHz LO, we have a phase shifter (delay line) already installed. So we can easily change the demodulation phase of REFL_2I.
Tonight, we switched the RF CARM signal to REFL_2I.
To do so, I changed the signal going to the REFL1 input of the common mode board from POX_1I to REFL_2I.
I moved a BNC-T installed at the output of POX_1I to the REFL_2I output to split the REFL_2I signal and send it to the CM board.
Since the gain of the REFL_2I was about 20dB lower than that of POX_1I, I increased the gain of the SR560, which is installed between the REFL_2 demodulation board and the CM board, from 1 to 10.

With some gain tweaks, we were able to hand off the CARM from REFL_DC to REFL_2I. We also succeeded in switching the REFL_2I ADC channel from PD11 to PD2_DC (the output of the length path from the CM board). This switching is necessary in order to engage the boost on the CM board.

There remains some offset in the CARM when the arm power is maximized. This is expected because the REFL_2I demodulation phase is probably not exactly right.
I will optimize the demodulation phase tomorrow.
  1530   Tue Apr 28 17:51:13 2009 robHowToLockingsetting the RF CARM demod phase

Quote:

To set the demod phase for RF CARM, sensed at REFL2 (REFL 166I), it suffices to set the demod phase for REFL2 to be the optimal phase for controlling SRCL in a no-arm state.

 

For POX33, the ideal phase for single arm locking does not yield a zero-offset CARM signal.  So the offset needs to be manipulated digitally. 

Attachment 1: XARM_phases_POX.pdf
XARM_phases_POX.pdf XARM_phases_POX.pdf XARM_phases_POX.pdf XARM_phases_POX.pdf
Attachment 2: CARM_phases_POX.pdf
CARM_phases_POX.pdf CARM_phases_POX.pdf CARM_phases_POX.pdf CARM_phases_POX.pdf
  1529   Tue Apr 28 16:36:24 2009 robHowToLockingsetting the RF CARM demod phase

To set the demod phase for RF CARM, sensed at REFL2 (REFL 166I), it suffices to set the demod phase for REFL2 to be the optimal phase for controlling SRCL in a no-arm state.

Attachment 1: CARM_phases_REFL.pdf
CARM_phases_REFL.pdf CARM_phases_REFL.pdf CARM_phases_REFL.pdf CARM_phases_REFL.pdf
Attachment 2: SRCL_phases_REFL.pdf
SRCL_phases_REFL.pdf SRCL_phases_REFL.pdf SRCL_phases_REFL.pdf SRCL_phases_REFL.pdf
ELOG V3.1.3-